OpenStudy (anonymous):

which points are collinear 3,1 , 5,5 and 7,9,

OpenStudy (anonymous):

don forget to show work

OpenStudy (angela210793):

5:5 cause they are on the line y=5

OpenStudy (anonymous):

how do you show the work though?????

OpenStudy (angela210793):

Hmm....I don't really know how to explain in any other way..sorry

OpenStudy (bahrom7893):

Jackie do the same thing as before..: Find the line that connects that (3,1) and (5,5).

OpenStudy (bahrom7893):

the equation is: y-y1 = M(x-x1) 5-1 = M(5-3) 4 = M * 2, M = 2

OpenStudy (anonymous):

well its cuz idk how to do that sorry -_-

OpenStudy (bahrom7893):

so now: y - y1 = 2 ( x - x1 ) y - 1 = 2( x - 3 ) y - 1 = 2x - 6 y = 2x - 5

OpenStudy (bahrom7893):

Now try the last point and see if it's on the line: (7,9) when x = 7 y = 2 * 7 - 5 = 14 - 5 = 9 so as you see: x = 7, y = 9 yea that's the same as our point: (7,9)

OpenStudy (bahrom7893):

so all 3 points are on the same line

OpenStudy (anonymous):

so there all collinear

OpenStudy (bahrom7893):

yea

OpenStudy (anonymous):

ok thank you